Solve for x in degrees, where 0 degrees <= x < 360 degrees . 2sin^2 x - sin x = 1


1
Expert's answer
2021-02-24T07:41:54-0500

Solution.

"2\sin^2{x}-\sin{x}=1, \text{ }0^{\degree}\leq x <360^{\degree}."

Make a replacement: "t=\sin{x}," then "2t^2-t-1=0."

Solve this equation:


"D=b^2-4ac=1-4\cdot 2\cdot (-1)=9,""t_1=\frac{-b-\sqrt{D}}{2a}=\frac{1-3}{2\cdot 2}=-\frac{1}{2};
t_2=\frac{-b+\sqrt{D}}{2a}=\frac{1+3}{2\cdot 2}=1."

Back to the replacement:

1)

"\sin{x}=-\frac{1}{2}, \newline
x=(-1)^n\arcsin{(-\frac{1}{2})}+\pi n, n \in Z, \newline
x=(-1)^{n+1}\arcsin{\frac{1}{2}}+\pi n, n \in Z, \newline

x=(-1)^{n+1}\frac{\pi}{6}+\pi n, n \in Z, \newline
x=(-1)^{n+1}30^{\degree}+180^{\degree} n, n \in Z."

At

"n=1: x=30^{\degree}+180^{\degree}=210^{\degree}, \newline
n=2: x=-30^{\degree}+180^{\degree}\cdot 2=330^{\degree}. \newline"

2)

"\sin{x}=1, \newline
x=\frac{\pi}{2}+2\pi k, k \in Z, \newline
x=90^{\degree}+360^{\degree}k, k \in Z. \newline"

At

"k=0: x=90^{\degree}."

Answer. "x=210^{\degree}, x=330^{\degree}, x=90^{\degree}."